OpenStudy (anonymous):

A box containing 12 cans of soda (mass 4.30 kg) is initially at rest on a horizontal surface. Then she is pushed 1.20 m straight by a trained dog who exerts a constant force equal to 36.0 N. module Use the work-energy theorem to find the final speed of the box is: a) there is no friction between the box and the surface; b) the coefficient of kinetic friction between the box and the surface is equal to 0.30.

OpenStudy (anonymous):

Work= Force x distance Work = (36.0N)(1.20 m) = (1/2) m v^2 in the absence of friction (b) Kinetic energy = (1./2) m v^2 = net force x distance = (36.0 N - 0.30 m g)(1.20m)

OpenStudy (anonymous):

@whpalmer4